OpenStudy (anonymous):

Convert the angle 30°37'11" to decimal degrees and round to the nearest hundredth of a degree.

OpenStudy (anonymous):

help me!!!!!!!!!!!!!!

OpenStudy (anonymous):

30°37'11" 11" = (11/60)/60 = .003055555555555555° 37' = 37/60 = .6166666666666666° 30° 30°+.6166666666666666°+ .003055555555555555° =30.6197°

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

your welcome